OpenStudy (anonymous):

Jessica is 2 times as old as Luis. Eighteen years ago, Jessica was 5 times as old as Luis. How old is Jessica now?

OpenStudy (haseeb96):

Too hard question it is shereen will help u she is smart in solving problems.hahha

OpenStudy (sleepyjess):

Do you know how to start this?

OpenStudy (sleepyjess):

@janxsgirl

OpenStudy (lichking):

this is linear equations damn easy

OpenStudy (sleepyjess):

To start yuo will need to use theses equations where J is Jessica and L is Luis. J=2L J-15=(L-15)*5

OpenStudy (sleepyjess):

From there it is pretty much guess and check. Lets start with the number 20 for Luis J=2(20) J=40 40-15=(20-15)*5 25=5*5 25=25 These equation are equal so Jessica is now 40

OpenStudy (anonymous):

40 is incorrect

OpenStudy (sleepyjess):

Oh whoops it was supposed to be 18. I was pretty sure it said 15 J=2L J-18=(L-18)*5 Now guess and check

OpenStudy (sleepyjess):

Can you get it from there or do you want another example?

OpenStudy (anonymous):

j=2l j-18=5(l-18) j-18=5(j/2-18) j-18=5/2j-90 3/2j=72 j=2/3*72 = 48 jessica is 48

OpenStudy (sleepyjess):

\(\Large\frak\color{lightgreen}{Correct}\)

OpenStudy (anonymous):

Grats KhanAcademy